OpenStudy (anonymous):

what is the value of the y variable in the soution to the following system of equations? 3x+y=14 -3x+27=1

OpenStudy (anonymous):

3x + y = 14 -3x + 27 = 1 --------------add wait....is there supposed to be a y term in the 2nd equation ? Is it 27y ??

OpenStudy (anonymous):

sorry the second equation is -3x+2y=1

OpenStudy (anonymous):

3x + y = 14 -3x + 2y = 1 ------------add 3y = 15 --- divide by 3 y = 15/3 y = 5 any questions ?

OpenStudy (anonymous):

you see that when adding the equations, the x terms cancel each other out. Leaving you to just solve for y